Snookers Lumber can convert logs into either lumber or plywood. In a given day, the mill turns out twice as many units of playwood as lumber. It makes a profit of $25 on a unit of lumber and $50 on a unit of plywood. How many of each unit must be produced and sold in order to make a profit of $13000?
Snooker Lumber must produce and sell Units of lumber and Units of plywood to make a profit of $13000.

Answer by mananth(12270)

lumber l units
plywood p units p=2l
25l+50p=13000
but p=2l
25l + 50*2l=13000
125l=13000
/125
l=13000/125
Lumber units 104
plywood units 104 *2=208 units